# Transport in integrable and perturbed easy-axis Heisenberg chain: Thouless approach

### v1

@@ Abstract
We study transport in the spin chains by employing the Thouless approach based on the level sensitivity to the boundary conditions, $R$. Although spin transport in the integrable easy-axis XXZ model is diffusive, corresponding $R$ is much closer to ballistic chains than to chaotic diffusive systems. In the case of the grand canonical ensemble this observation can be rigorously justified, while in the case of the canonical ensemble it can be demonstrated by numerical calculations. Integrability breaking perturbation (IBP) strongly reduces $R$ which reveals a pronounced minimum at the crossover from anomalous diffusive to normal dissipative transport. This minimum coincides with the onset of the universality of the random matrix theory. Results for various IBP suggest a discontinuous jump of the spin conductivity in the thermodynamic limit, and moreover that its value universally decreases when the strength of IBP decreases.
Introduction. The transport and other properties of integrable quantum lattice models, the prominent example being the one-dimensional, anisotropic Heisenberg XXZ spin-$1/2$ chain, have in last decades reached higher level of understanding (for recent review see e.g. [1]) through the extended concepts of local [2] and quasilocal conserved quantities (CQ) [3, 4], via the analytical results emerging from the generalized hydrodynamics (GHD) [5, 6, 7, 8, 9] and with the application of various powerful numerical methods [10, 11, 12, 13, 14, 15]. While the standard argument is that mastering of integrable models will be the crucial step to the proper description of closely related perturbed models, the advance in this direction is so far modest. In the case of the Heisenberg-type spin chains it appears evident that the integrability-breaking perturbations (IBP) suppress the finite-temperature ballistic transport, characterized by the spin stiffness $D(T>0)>0$, turning it into the normal dissipative one [16, 17, 18, 19, 20]. Still, analytical approaches and tools dealing with such perturbed systems are so far quite restricted [21, 22, 23] .
@@
The level sensitivity distinguishes also between ballistic and normal diffusive transport. In the former case it grows (up to logarithmic corrections) as the inverse of the mean level spacing, $R\propto 1/\Delta\epsilon$ whereas in generic diffusive systems this dependence is much weaker, namely $R\propto 1/\sqrt{\Delta\epsilon}$. We show that the diffusive transport in the easy-axis integrable XXZ chain ($\Delta>1$) is anomalous in this respect. Namely, diffusive integrable chains reveal $R\propto 1/\Delta\epsilon$ in the grandcanonical ensemble (GCE), as well as in the canonical ensemble (CE) at total $S^{z}_{tot}=0$. Therefore, the level sensitivity of diffusive integrable system differs from $R$ in the ballistic systems only by logarithmic corrections, $\log(\Delta\epsilon)$, and is much larger than in generic diffusive models.
Large values of the level sensitivity in GCE, $R_{\text{GCE}}\gg 1$, can be explained by finding its strict lower bound through the overlap with the product of simplest integrals of motion of the XXZ model. On the other hand, level sensitivity obtained in CE, $R_{\text{CE}}$, strongly decreases with $\Delta$. In a particular limit $\Delta\to\infty$, representing the ’folded’ XXZ model [34, 35, 30], we even find strictly $R_{\text{CE}}=0$. This result strongly contrasts with $R_{\text{GCE}}$ for which the lower bound increases with $\Delta$.
We also observe that $R_{\text{CE}}$ in finite systems exhibits a nonmonotonic dependence on the strength of IBP, $g$. The minimum at $g=g^{*}(L)$ coincides with the crossover where the finite-size system starts to comply with the random matrix theory (RMT) [36, 37]. Therefore, the minimum of $R$ allows us to delimit two regimes: $g>g^{*}(L)$ when the system behaves as normal dissipative and $g<g^{*}(L)$ when the system properties are still dominated by finite-size effects. The meaning of the crossover can be confirmed also with standard indicators of RMT. Moreover, taking into account the restriction $g>g^{*}(L)$, one can then follow in a controlled way the variation with $g$ of the d.c. spin conductivity in CE, $\sigma_{\text{CE}}$. Our results for two different classes of IBP seem to support a universal limit, i.e., vanishing of $\sigma_{\text{CE}}(g\to 0)$ in the thermodynamic limit.
XXZ Model with flux. We consider the XXZ spin ($s=1/2$) chain of length $L$ with periodic boundary conditions (PBC), introducing also the effect of finite flux, $\varphi$, into the exchange term
$H=J\sum_{l}\left[\frac{1}{2}(\mathrm{e}^{i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c.})+\Delta S^{z}_{l+1}S^{z}_{l}\right]+gH^{\prime}.$ (1)
We consider the easy-axis regime with the anisotropy $\Delta>1$ and set $J=1$ as the unit of energy. The spin model can be mapped on the model of interacting spinless fermions, where $\varphi L$ represents a magnetic flux through the ring [31, 32, 15]. We analyze two different prototype forms of IBP. In the main text, we discuss the next-nearest neighbor (nnn) interaction with $g=\Delta_{2}$ and $H^{\prime}=\sum_{l}S^{z}_{l+2}S^{z}_{l}$, whereas in the Supplement [38] we study also the nnn exchange with $g=J_{2}$ and $H^{\prime}=\sum_{l}[\mathrm{e}^{2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}]$ which maps to a nnn hopping in the fermionic chain.
@@
Level sensitivity and spin transport. The gauge transformation $S^{\pm}_{l}\to e^{\mp il\varphi}S^{\pm}_{l}$, $S^{z}_{l}\to S^{z}_{l}$ applied to the Hamiltonian (1) eliminates the flux dependence on all sites except the boundary term that takes the form $J/2(S^{+}_{1}S^{-}_{L}e^{i\varphi L}+{\rm H.c.})$. Then, the energy spectra, $H|n\rangle=\epsilon_{n}(\varphi)|n\rangle$, obtained for $\varphi=0$ and $\varphi=\delta\varphi=\pi/L$ correspond, respectively, to periodic and antiperiodic ($J\to-J$) boundary conditions. Consequently, the level sensitivity to the boundary conditions is related to the derivative $\epsilon^{\prime}_{n}(\varphi)$ that can be further simplified using an identity for the spin-current operator, $j={\rm d}H/{\rm d}\phi$. Following the reasoning from Refs. [31, 39] one finds
@@
Here, overline means averaging over the eigenstates (in the relevant part of the spectrum) and $\Delta\epsilon$ is the average level spacing. We note that integrable XXZ model in the easy-plane regime $\Delta<1$ exhibits a ballistic transport. It is quantified by the spin stiffness (Drude weight) [2], $D=\overline{j^{2}_{nn}}/L>0$ that is finite in the thermodynamics limit $L\to\infty$. Consequently, in ballistic models one obtains $R=\pi\sqrt{D}/(\sqrt{L}\Delta\epsilon)$.
In the case of the easy-axis regime, $\Delta>1$, the ballistic component vanishes in GCE and in CE with $S^{z}_{tot}=0$. In both cases one finds $D=0$ for $L\to\infty$. However in the GCE, one may still find a simple bound on the level sensitivity. To this end we employ the Mazur bound [40, 2], $\overline{j^{2}_{nn}}\geq[{\rm Tr}(Qj)]^{2}/{\rm Tr}(Q^{2})$ where $Q$ is a conserved operator. We choose $Q$ as a product of two simplest local conserved operators, $Q=Q_{1}Q_{3}$, where $Q_{1}=\sum_{l}S^{z}_{l}$ is the total magnetization and $Q_{3}$ is the energy current,
@@
$\overline{j^{2}_{nn}}\geq\frac{\left[{\rm Tr}\left(Q_{1}Q_{3}j\right)\right]^{2}}{{\rm Tr}\left(Q^{2}_{1}Q^{2}_{3}\right)}=\frac{J^{2}\Delta^{2}L}{2J^{2}(L-2)+4L\Delta^{2}},$ (5)
@@
The above bound is inapplicable in the CE with $S^{z}_{tot}=0$ because $Q_{1}=0$. Fig. 1(b) shows finite-size results for $\overline{j^{2}_{nn}}$. Here and further on, they are obtained using exact diagonalization (ED). For $L\leq 22$, we carry out full diagonalization for sectors with the wavevectors $0<q<\pi$ and use half of the states from the middle of the spectra. For the largest $L=24$ we employ the shift-invert ED and randomly sample over $N\sim 150$ eigenstates from the middle of the spectra, in each sector with $0<q<\pi$ (see Supplement [38] for details).
One observes that $\overline{j^{2}_{nn}}$ is well below the GCE bound and that the difference between GCE and CE results increases with $L$ and $\Delta$. In order to elucidate such $\Delta$-dependence, we have numerically studied also the ’folded’ XXZ model [34, 35, 30] that represents $\Delta\to\infty$ limit of the XXZ chains, $H_{f}=(J/2)\sum_{l}P_{l}\left[\mathrm{e}^{i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c}\right]$ with the projection operator $P_{l}=(S^{z}_{l+2}+S^{z}_{l-1})^{2}$ that introduces conservation of the number of domain walls. In this model we find (so far numerically) strictly $\overline{j^{2}_{nn}}=0$ for any $L$. Consequently, one gets $R_{\text{CE}}=0$ while $R_{\text{GCE}}>0$ as it follows from the strict bound in Eq. (5).
Following results in Fig. 1(b) we present in Fig. 2(a) the dependence of $R_{\text{CE}}$ on the average level spacing $\Delta\epsilon$. Results show that for finite $\Delta<\infty$, the anomalous dependence $R\propto 1/\Delta\epsilon$ holds true up to logarithmic corrections, ${\cal O}[\log(\Delta\epsilon)]$, being thus close to ballistic systems rather than to generic diffusive ones, as discussed further on.
@@
where $N_{st}$ is the Hilbert-space dimension, and $\overline{|j_{mn}|^{2}}$ are averaged matrix elements for small $|\epsilon_{m}-\epsilon_{n}|$. Upon introducing a flux $\varphi\neq 0$, the degeneracy of energy levels persists only in momentum sectors $q=0$ and $q=\pi$. All other sectors obey the RMT relations [43], which in our system follow the Gaussian orthogonal ensemble (GOE) universality, requiring $Y=\overline{|j_{mn}|^{2}}/\overline{j_{nn}^{2}}=1/2$. Therefore, RMT establishes a link between the level sensitivity, Eq. (3) and the spin conductivity, Eq. (6),
@@
Consequently, introducing of IBP must be accompanied by a substantial drop of the level sensitivity from $R\propto 1/(L\Delta\epsilon)$ in the integrable chain (either ballistic or diffusive) to $R\propto 1/\sqrt{L\Delta\epsilon}$ in the chaotic system. In the GCE, such scenario is a direct consequence of the bound in Eq. (5). In the CE with finite $\Delta<\infty$ we confirm this scenario via numerical calculations and all result henceforth apply to the latter.
Figure 2: Results for CE with $S^{z}_{tot}=0$ from full ED (filled symbols) and shift-invert method (open symbols). (a) Scaled level sensitivity $LR_{\text{CE}}$ vs $1/\Delta\epsilon$, where $\Delta\epsilon$ is average level spacing in sectors with fixed momenta. Dashed lines are guidelines showing $1/\Delta\epsilon$ and $1/\sqrt{\Delta\epsilon}$ dependence, respectively. (b)-(d) $R_{\text{CE}}$ vs. IBP strength $\Delta_{2}$ for $\Delta=1.5,2.0,3.0$. Thick black lines mark the minima of $R(\Delta_{2})$ while regions with $R<1$ are shaded.
@@
Fig. 2(a) confirms that substantial IBP ($g=\Delta_{2}$) causes qualitative changes of $R_{\text{CE}}$, as discussed in the preceding paragraphs. More systematic dependence of $R_{\text{CE}}$ on the IBP strength $\Delta_{2}$ is presented in Figs. 2(b)- 2(d) for different $\Delta>1$. One observes that $R_{\text{CE}}$ develops a clear minimum at finite $\Delta_{2}$ that drifts towards weaker IBP when $L$ increases. It should be noted, that we are interested in the regime $R_{\text{CE}}\gtrsim 1$. For $R_{\text{CE}}\lesssim 1$ (shaded areas in Fig. 2) the system is too small, i.e. the level spacing is too large, so that changing of the flux causes neither level crossings nor avoided crossings (within the same $S^{z}_{tot}$ and momentum sector).
@@
We note that results for $\sigma_{\text{CE}}$ obtained from various methods are consistent provided that $g>g^{*}(L)$. One also observes that $\sigma_{\text{CE}}$ strongly depends on $\Delta$ resembling the analogous dependence of $R_{\text{CE}}$ on $\Delta$ shown Fig. 2(a). Apparently for $g>g^{*}(L)$, the spin conductivity follows the relation $\sigma_{\text{CE}}\propto|g|$. Here we can make a link with the diffusion constant, ${\cal D}$. In generic dissipative system it is related to $\sigma$ via the Einstein relation, that at large $T$ reads ${\cal D}=4\sigma$. The diffusion has been studied extensively in the integrable case and evaluated mostly in the GCE, both analytically within the GHD [25] and numerically via various methods [26, 11, 14, 13, 15] leading to values ${\cal D}_{\text{GCE}}>0.4$ in the regime $\Delta>1$ Evidently, $\sigma_{\text{CE}}$ shown in Fig. 4 for perturbation $g>g^{*}(L)$ is well below the latter. This discrepancy suggests a discontinuous change of ${\cal D}$ upon the introduction of IBP in agreement with Refs. [30, 15]. Here, the underlying assumptions are the Einstein relation and the equivalence of CE and GCE. Numerical testing of the latter equivalence for weakly perturbed systems is extremely demanding since the difference between both ensembles approximately follows the relation ${\cal D}_{\text{GCE}}-{\cal D}_{\text{CE}}\propto 1/(g^{2}L)$ [15].
The relevant question is how general are conclusions presented above for a particular form of IBP with $g=\Delta_{2}$. In the Supplemental Material [38] we therefore present analogous results for another class of IBP, i.e. for $g=J_{2}$ representing the nnn exchange. Conclusions are quite similar to the above ones for $g=\Delta_{2}$, even that concerning the apparent vanishing $\sigma_{\text{CE}}(g\to 0)\to 0$. The difference is that $J_{2}>0$ introduces nonvanishing matrix elements $\overline{|j_{nn}|^{2}}>0$ also in the folded model.
Conclusions and Discussion We have studied sensitivity of the energy levels to the flux (or equivalently twisted boundary conditions), $R$, and related transport quantities, within the XXZ chain in the easy-axis regime. Although the transport in the integrable model is diffusive, obtained $R$ is very close to one in integrable ballistic systems and, up to logarithmic corrections, exhibits the same dependence on the level spacing, $R\propto 1/\Delta\epsilon$. It strongly contrasts with chaotic diffusive systems where $R\propto 1/\sqrt{\Delta\epsilon}$. In the case of GCE , we have found for integrable chain a strict bound $R_{\text{GCE}}\geq 0.4\pi/(L\Delta\epsilon)$ that holds for arbitrary anisotropy $\Delta\geq 1$ and all system sizes, $L$. In the CE and for finite anisotropy $\Delta<\infty$, the level sensitivity still varies as $R_{\text{CE}}\propto 1/\Delta\epsilon$, however $R_{\text{CE}}$ is much smaller than $R_{\text{GCE}}$. The difference between $R_{\text{GCE}}$ and $R_{\text{CE}}$ grows with the systems sizes and anisotropy. In the limiting case $\Delta\to\infty$ we have found that $R_{\text{CE}}$ strictly vanishes for all system sizes whereas $R_{\text{GCE}}$ is bounded from below.
@@
In the chaotic regime, RMT establishes the Thouless-like relation between $R$ and the d.c. spin conductivity, $\sigma$. Using this relation we have numerically studied $\sigma$ in CE and found for $g>g^{*}(L)$ that $\sigma_{CE}\propto|g|$. For weak perturbations, $\sigma_{CE}$ becomes much smaller than $\sigma_{GCE}$ estimated from the diffusion constant for integrable systems. Assuming that the Einstein relation and the equivalence between CE and GCE hold true for all model parameters, our results support a discontinuous variation of $\sigma_{GCE}=\sigma_{CE}$ with IBP. It should be acknowledged that the limit of a weak perturbation is highly nontrivial since the difference between CE and GCE results goes as, $1/(g^{2}L)$ [15]. Consequently, the convergence of results in both ensembles would be hard to establish numerically.
@@
Transport in integrable and perturbed easy-axis Heisenberg chain: Thouless approach
@@
In the Supplemental Material we present additional results with next-nearest-neighbor exchange as the perturbation and details about shift-invert calculations.
@@
$H^{\prime}=\sum_{l}[\mathrm{e}^{2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}].$ (S1)
@@
Figure S1: Level sensitivity $R_{CE}$ vs. perturbation strength $J_{2}$ within sector $S^{z}_{tot}=0$, as calculated via full ED in systems with $L=18-22$ sites, for: (a) $\Delta=1.5$, (b) $\Delta=2.0$, and (c) $\Delta=3.0$.

### v2

# Thouless approach in transport in integrable and perturbed easy-axis Heisenberg chains

@@ Abstract
We study transport in spin chains employing the Thouless approach based on the level sensitivity to the boundary conditions, $R$. Although spin transport in the integrable easy-axis XXZ model is diffusive, corresponding $R$ is much closer to ballistic chains than to chaotic diffusive systems. In the case of the grand canonical ensemble, this observation can be rigorously justified, while in the case of the canonical ensemble it can be demonstrated by numerical calculations. Integrability breaking perturbation (IBP) strongly reduces $R$ which reveals a pronounced minimum at the crossover from anomalous diffusive to normal dissipative transport. This minimum coincides with the onset of the universality of the random matrix theory. Results for various IBP suggest a discontinuous jump of the spin conductivity in the thermodynamic limit, and moreover that its value is proportional to the strength of the IBP.
Introduction. The properties of integrable quantum lattice models, the prominent example being the one-dimensional, anisotropic Heisenberg XXZ spin-$1/2$ chain, have in last decades reached higher level of understanding (for recent review see e.g. [1]) through the extended concepts of local [2] and quasilocal conserved quantities (CQ) [3, 4], via the analytical results emerging from the generalized hydrodynamics (GHD) [5, 6, 7, 8, 9] and with the application of various powerful numerical methods [10, 11, 12, 13, 14, 15]. While the standard argument is that mastering integrable models will be the crucial step to the proper description of closely related perturbed models, the advance in this direction is so far modest. In the case of the Heisenberg-type spin chains it appears evident that the integrability-breaking perturbations (IBP) suppress the finite-temperature ballistic transport, characterized by the spin stiffness $D(T>0)>0$, turning it into the normal dissipative one [16, 17, 18, 19, 20]. Still, analytical approaches and tools dealing with such perturbed systems are so far quite restricted [21, 22, 23] .
@@
The level sensitivity distinguishes between ballistic and normal diffusive transport. In the former case it grows (up to logarithmic corrections) as the inverse of the mean level spacing, $R\propto 1/\Delta\epsilon$ whereas in generic diffusive systems this dependence is much weaker, namely $R\propto 1/\sqrt{\Delta\epsilon}$. We show that the diffusive transport in the easy-axis integrable XXZ chain ($\Delta>1$) is anomalous in this respect. Namely, diffusive integrable chains reveal $R\propto 1/\Delta\epsilon$ in the grandcanonical ensemble (GCE), as well as in the canonical ensemble (CE) at total $S^{z}_{tot}=0$. Therefore, the level sensitivity of diffusive integrable system differs from $R$ in the ballistic systems only by logarithmic corrections, $\log(\Delta\epsilon)$, and is much larger than in generic diffusive models. Large values of the level sensitivity in GCE, $R_{\text{GCE}}\gg 1$, are explained by strict lower bound.
We also observe that level sensitivity in CE, $R_{\text{CE}}$, obtained for finite systems exhibits a nonmonotonic dependence on the strength of IBP, $g$. The minimum at $g=g^{*}(L)$ coincides with the crossover where the finite-size system starts to comply with the random matrix theory (RMT) [34, 35]. Therefore, the minimum of $R$ allows us to delimit two regimes: $g>g^{*}(L)$ when the system behaves as normal dissipative and $g<g^{*}(L)$ when the system properties are still dominated by finite-size effects. Taking into account the restriction, $g>g^{*}(L)$, one can then follow in a controlled way the variation with $g$ of the d.c. spin conductivity in CE, $\sigma_{\text{CE}}$. Our results for two classes of IBP seem to support a universal limit, i.e., vanishing of $\sigma_{\text{CE}}(g\to 0)$ in the thermodynamic limit.
XXZ Model with flux. We consider the XXZ spin ($s=1/2$) chain of length $L$ with periodic boundary conditions (PBC), introducing also a flux, $\varphi$,
$H=J\sum_{l}\left[\frac{1}{2}(\mathrm{e}^{-i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c.})+\Delta S^{z}_{l+1}S^{z}_{l}\right]+gH^{\prime}.$ (1)
We consider the easy-axis regime with $\Delta>1$ and set $J=1$. The spin model can be mapped on the model of interacting spinless fermions, where $\varphi L$ represents a magnetic flux through the ring [31, 32, 15]. We analyze two forms of IBP. In the main text, we discuss the next-nearest neighbor (nnn) interaction with $g=\Delta_{2}$ and $H^{\prime}=\sum_{l}S^{z}_{l+2}S^{z}_{l}$, whereas in the Supplemental Material [36] we study the nnn exchange with $g=J_{2}$ and $H^{\prime}=\sum_{l}[\mathrm{e}^{-2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}]$ which maps to a nnn hopping in the fermionic chain.
@@
Level sensitivity and spin transport. The gauge transformation $S^{\pm}_{l}\to e^{\pm il\varphi}S^{\pm}_{l}$, $S^{z}_{l}\to S^{z}_{l}$ applied to the Hamiltonian (1) eliminates the flux dependence on all sites except the boundary term that takes the form $J/2(S^{+}_{1}S^{-}_{L}e^{-i\varphi L}+{\rm H.c.})$. Then, the energy spectra, $H|n\rangle=\epsilon_{n}(\varphi)|n\rangle$, obtained for $\varphi=0$ and $\varphi=\delta\varphi=\pi/L$ correspond, respectively, to periodic and antiperiodic ($J\to-J$) boundary conditions. Consequently, the level sensitivity to the boundary conditions is related to the derivative $\epsilon^{\prime}_{n}(\varphi)$ that can be further simplified using an identity for the spin-current operator, $j={\rm d}H/{\rm d}\varphi$. Following the reasoning from Refs. [31, 37] one finds
@@
Here, overline means averaging over the eigenstates (in the relevant part of the spectrum) and $\Delta\epsilon$ is the average level spacing. We note that integrable XXZ model with $\Delta<1$ exhibits a ballistic transport. It is quantified by the spin stiffness (Drude weight) [2], $D=\overline{j^{2}_{nn}}/L>0$ that is finite in the thermodynamics limit $L\to\infty$. Consequently, in ballistic models one obtains $R=\pi\sqrt{D}/(\sqrt{L}\Delta\epsilon)$.
In the case of the easy-axis regime, $\Delta>1$, the ballistic component vanishes in GCE and in CE with $S^{z}_{tot}=0$. In both cases one finds $D=0$ for $L\to\infty$. However in the GCE, one may still find a simple bound on the level sensitivity. To this end we employ the Mazur bound [38, 2], $\overline{j^{2}_{nn}}\geq\langle jQ\rangle^{2}/\langle Q^{2}\rangle$, where $Q$ is a conserved operator, $\langle...\rangle={\rm Tr}(...)/N_{st}$ and $N_{st}$ is the Hilbert-space dimension. We choose $Q$ as a product of two simplest local conserved operators, $Q=Q_{1}Q_{3}$, where $Q_{1}=\sum_{l}S^{z}_{l}$ is the total magnetization and $Q_{3}$ is the energy current,
@@
$\overline{j^{2}_{nn}}\geq\frac{\langle jQ_{1}Q_{3}\rangle^{2}}{\langle Q^{2}_{1}Q^{2}_{3}\rangle}=\frac{J^{2}\Delta^{2}L}{2J^{2}(L-2)+4L\Delta^{2}},$ (5)
@@
The above bound is inapplicable in the CE with $S^{z}_{tot}=0$ because $Q_{1}=0$. Fig. 1(b) shows finite-size results for $\overline{j^{2}_{nn}}$. Here and further on, they are obtained using exact diagonalization (ED). For $L\leq 22$, we carry out full diagonalization for sectors with the wavevectors $0<q<\pi$ and use half of the states from the middle of the spectra. For the largest $L=24$ we employ the shift-invert ED and randomly sample over $N\sim 150$ eigenstates from the middle of the spectra, in each sector with $0<q<\pi$ (see Supplemental Material [36] for details).
One observes that $\overline{j^{2}_{nn}}$ is well below the GCE bound and that the difference between GCE and CE results increases with $L$ and $\Delta$. In order to elucidate such $\Delta$-dependence, we have numerically studied also the ’folded’ XXZ model [39, 40, 30] that represents $\Delta\to\infty$ limit of the XXZ chains, $H_{f}=(J/2)\sum_{l}P_{l}\left[\mathrm{e}^{-i\varphi}S^{+}_{l+1}S^{-}_{l}+\mathrm{H.c}\right]$ with the projection operator $P_{l}=(S^{z}_{l+2}+S^{z}_{l-1})^{2}$ that introduces conservation of the number of domain walls. In this model we find (so far numerically) strictly $\overline{j^{2}_{nn}}=0$ for any $L$. Consequently, one gets $R_{\text{CE}}=0$ while $R_{\text{GCE}}>0$. The latter follows from the strict bound in Eq. (5) and originates from the ballistic transport in all sectors with $S^{z}_{tot}\neq 0$. Analytical support for vanishing $R_{\text{CE}}$ is discussed in the Supplemental Material. [36] Following results in Fig. 1(b) we present in Fig. 2(a) the dependence of $R_{\text{CE}}$ on the average level spacing $\Delta\epsilon$. Results show that for finite $\Delta<\infty$, the anomalous dependence $R\propto 1/\Delta\epsilon$ indeed holds true, being thus close to ballistic systems rather than to generic diffusive ones.
@@
where $\overline{|j_{mn}|^{2}}$ are averaged matrix elements for small $|\epsilon_{m}-\epsilon_{n}|$. Upon introducing a flux $\varphi\neq 0$, the degeneracy of energy levels persists only in momentum sectors $q=0$ and $q=\pi$. All other sectors obey the RMT relations [43], which in our system follow the Gaussian orthogonal ensemble (GOE) universality, requiring $Y=\overline{|j_{mn}|^{2}}/\overline{j_{nn}^{2}}=1/2$. Therefore, RMT establishes a link between the level sensitivity, Eq. (3) and the spin conductivity, Eq. (6),
@@
Eq. (7) means that the level sensitivity in chaotic systems scales with the system size as $R\propto 1/\sqrt{L\Delta\epsilon}$. It is a general (i.e. model independent) property that follows solely from RMT and the finiteness of $\sigma$. In the case of integrable systems, $R$ should follow a simple bound $R\geq c/(L^{m+1}\Delta\epsilon)$ where $c$ is a constant of order one. Independently of the studied model, a sufficient condition for the latter to hold is that the current operator has a non-vanishing projection on products of a finite number of LIOMs, $\langle jQ_{1}Q_{2}\ldots Q_{m}\rangle\neq 0$, see the discussion in the Supplemental Material [36]. We stress that the latter projection implies ballistic transport only for $m=1$. We also recall that the level spacing, $\Delta\epsilon$, decays exponentially with $L$. Consequently, introducing of IBP should in general be accompanied by a substantial drop of the level sensitivity from $R\propto 1/\Delta\epsilon$ in the integrable chain (either ballistic or diffusive) to $R\propto 1/\sqrt{\Delta\epsilon}$ in the chaotic system. All other terms which enter $R$ can be bounded by a power-law function $\sim 1/L^{m+1}$ so that they appear only as logarithmic corrections, $\sim\log(\Delta\epsilon)$.
Figure 2: Results for CE with $S^{z}_{tot}=0$ from full ED (filled symbols) and shift-invert method (open symbols). (a) Scaled level sensitivity $LR_{\text{CE}}$ vs $1/\Delta\epsilon$, where $\Delta\epsilon$ is average level spacing in sectors with fixed momenta. Dashed lines are guidelines showing $1/\Delta\epsilon$ and $1/\sqrt{\Delta\epsilon}$ dependence, respectively. (b)-(d) $R_{\text{CE}}$ vs. IBP strength $\Delta_{2}$ for $\Delta=1.5,2.0,3.0$. Regions with $R<1$ are shaded, whereas black triangles correspond to minima of $R(\Delta_{2})$, approximated from a quadratic interpolation of three subsequent data points.
@@
All result henceforth apply to CE. Figure 2(a) confirms that $R_{\text{CE}}$ in the integrable XXZ model shows the dependence $R_{\text{CE}}\propto 1/\Delta\epsilon$ for all finite $\Delta$. Comparing this result with $R_{\text{CE}}=0$ in the ’folded’ model, one finds that the limits $L\to\infty$ and $\Delta\to\infty$ do not commute. The limit $\Delta\to\infty$ is taken first in the ’folded’ model. Fig. 2(a) confirms also that substantial IBP ($g=\Delta_{2}$) causes qualitative changes of $R_{\text{CE}}$, as discussed in the preceding paragraphs. More systematic dependence of $R_{\text{CE}}$ on the IBP strength $\Delta_{2}$ is presented in Figs. 2(b)- 2(d) for different $\Delta>1$. One observes that $R_{\text{CE}}$ develops a clear minimum at finite $\Delta_{2}$ that drifts towards weaker IBP when $L$ increases. It should be noted, that we are interested in the regime $R_{\text{CE}}\gtrsim 1$. For $R_{\text{CE}}\lesssim 1$ (shaded areas in Fig. 2) the system is too small, i.e. the level spacing is too large, so that changing of the flux causes neither level crossings nor avoided crossings (within the same $S^{z}_{tot}$ and momentum sector).
@@
We note that results for $\sigma_{\text{CE}}$ obtained from various methods are consistent provided that $g>g^{*}(L)$. One also observes that $\sigma_{\text{CE}}$ strongly depends on $\Delta$ resembling the analogous dependence of $R_{\text{CE}}$ on $\Delta$ shown Fig. 2(a). Apparently for $g>g^{*}(L)$, the spin conductivity follows the relation $\sigma_{\text{CE}}\propto|g|$. Here we can make a link with the diffusion constant, ${\cal D}$. In generic dissipative system it is related to $\sigma$ via the Einstein relation, that at large $T$ reads ${\cal D}=4\sigma$. The diffusion has been studied extensively in the integrable case and evaluated mostly in the GCE, both analytically within the GHD [25] and numerically via various methods [26, 11, 14, 13, 15] leading to values ${\cal D}_{\text{GCE}}>0.4$ in the regime $\Delta>1$ Evidently, $\sigma_{\text{CE}}$ shown in Fig. 4 for perturbation $g>g^{*}(L)$ is well below the latter. This discrepancy suggests a discontinuous change of ${\cal D}$ upon the introduction of IBP in agreement with Refs. [30, 15].
Conclusions and Discussion We have studied sensitivity of the energy levels to the flux (or equivalently twisted boundary conditions), $R$, and related transport quantities, within the XXZ chain in the easy-axis regime. Although the transport in the integrable model is diffusive, obtained $R$ is very close to one in integrable ballistic systems and, up to logarithmic corrections, exhibits the same dependence on the level spacing, $R\propto 1/\Delta\epsilon$. It strongly contrasts with chaotic diffusive systems where $R\propto 1/\sqrt{\Delta\epsilon}$. In the case of GCE , we have found for integrable chain a strict bound $R_{\text{GCE}}\geq 0.4\pi/(L\Delta\epsilon)$ that holds for arbitrary anisotropy $\Delta\geq 1$ and all system sizes, $L$. In the CE and for finite anisotropy $\Delta<\infty$, the level sensitivity still varies as $R_{\text{CE}}\propto 1/\Delta\epsilon$, however $R_{\text{CE}}$ is much smaller than $R_{\text{GCE}}$. The difference between $R_{\text{GCE}}$ and $R_{\text{CE}}$ grows with the systems sizes and anisotropy. In the limiting case $\Delta\to\infty$ we have found that $R_{\text{CE}}$ strictly vanishes for all system sizes whereas $R_{\text{GCE}}$ is bounded from below. In the Supplemental Material [36], we discuss results for nonzero magnetization sectors.
@@
In the chaotic regime, RMT establishes the Thouless-like relation between $R$ and the d.c. spin conductivity, $\sigma$. Using this relation we have numerically studied $\sigma$ in CE and found for $g>g^{*}(L)$ that $\sigma_{CE}\propto|g|$. For weak perturbations, $\sigma_{CE}$ becomes much smaller than $\sigma_{GCE}$ estimated from the diffusion constant for integrable systems. Assuming that the Einstein relation and the equivalence between CE and GCE hold true for all model parameters, our results support a discontinuous variation of $\sigma_{GCE}=\sigma_{CE}$ with IBP. It should be acknowledged that the limit of a weak perturbation is highly nontrivial since the difference between CE and GCE results goes as, $1/(g^{2}L)$ [15]. Consequently, the convergence of results in both ensembles is hard to establish numerically.
@@
Thouless approach in transport in integrable and perturbed easy-axis Heisenberg chains
@@
In the Supplemental Material we present additional results with next-nearest-neighbor exchange as the perturbation, details of shift-invert calculations, and extra considerations of the ’folded’ XXZ model, which can serve as the toy model for the behavior of the easy-axis $\Delta>1$ regime. We also show an analytical argument bounding the diagonal matrix elements of the spin current. Finally, we discuss the level sensitivity for nonzero magnetization sectors.
@@
$H^{\prime}=\sum_{l}[\mathrm{e}^{-2i\varphi}S^{+}_{l+1}S^{z}_{l}S^{-}_{l-1}+\mathrm{H.c.}].$ (S1)
@@
Figure S1: Level sensitivity $R_{CE}$ vs. perturbation strength $J_{2}$ within sector $S^{z}_{tot}=0$, as calculated via full ED in systems with $L=18-22$ sites, for: (a) $\Delta=1.5$, (b) $\Delta=2.0$, and (c) $\Delta=3.0$. Black triangles correspond to minima of $R(\Delta_{2})$, approximated from a quadratic interpolation of three subsequent data points.